A 5.0 V battery storing 43.0 kJ of energy supplies 1.5 A of current to a circuit. How much energy does the battery have left after powering the circuit for 1.0 h?
Answer #1

For 1 hour supply

power supplied battery = V*I = 5*1.5 = 7.5W

Energy = power*time = 7.5 * 3600 J

battery has intially 43.00 kJ of energy

so energy left = 43000 - 7.5*3600

                   = 16000

                    = 16 kJ